I abolutely cannot figure this out!
Solve 18t/7=2t-2/3
Can anyone please help?

Answer by josmiceli(19441)

get the terms with the unknown on the left side by subtracting 2t
from both sides
18t/7 - 2t = -2/3
now multipy both sides by 7
18t - 14t = -14/3
4t = -14/3
divide both sides by 4
t = -14/12 = -7/6